Campus Technology Services

Taking Steps for the Future

TechNews - May 2013

This is a summary of the activities supporting faculty, staff and student technology services at SUNY Oswego that occurred during April 2013.

Reminders - Please make sure that your co-workers know about these

  1. Routine Service Outage
    As part of Campus Technology Services' efforts to keep servers running at optimum levels, a regular maintenance period has been designated. From 6AM to 10AM on the second Sunday of each month, several services such as E-mail, editing of both departmental and personal web pages will not be available. The next scheduled outage is Sunday, June 9th, 2013.
  2. NYALERT
    For information on NYAlert: http://www.oswego.edu/administration/public_affairs/emergency/nyalert.html or 
    visit http://www.suny.edu/SUNYALERTsec to sign up.
  3. RAVE Guardian
    For information on Rave Guardian http://oswego.edu/student/blogs/?cat=11 or to sign up: https://www.getrave.com/login/oswego
  4. Patch Wednesdays
    CTS now has the capability of updating your campus owned computers for security fixes as well as functionality on a regularly scheduled basis. This may include software updates to programs such as Adobe Reader or Microsoft Word or to change Windows settings to better support your computers. In order to streamline this process we will only perform these updates on Wednesdays.  These changes will be implemented only after thorough internal CTS testing. You may possibly notice a slight slowdown in computer performance at that time. Updates are added to a CTS calendar so you may see what changes have occurred to your computer on any given day or that are scheduled to be done.
  5. Workshops calendar 
    Look here for a listing of all professional development workshop registrations, calendars and information. 
  6. CTS News
    Make sure to visit our CTS New blog for some of our latest information on updated or new services, future services, technology tips, and more.
  7. CTS Projects
    Check our CTS Projects for brief descriptions of the projects CTS is currently working on. 

Technology Support Center

TSC Staff have been working on implementing the new Service Now ticketing system.  We will be replacing the HEAT ticketing system with the more robust, easy to use, ITIL-based ServiceNow product. All trainings with staff have been complted and we will be going live on Friday, May 3.  One of the goals with implementing ServiceNow is to increase communication with the people we are working with to provide service to.    In order to achieve this, you will begin to receive email notifications.  These emails will be from the "Technology Support Center" from an oswego.service-now.com email address.  They will be sent upon the opening, servicing and resolving of issues.

We will be searching for a new Help Desk Support Specialist full-time position.  HR is working on advertising the position.

We completed our student recruitment efforts and have hired 4 Help Desk Assistants for the 2013-2014 academic year.

You can reach us by phone at 312.3456 or on site at 26 Lanigan Hall, on Facebook at "SUNY Oswego - Campus Technology Services" on Twitter at www.twitter.com/sunyoswegocts or our website.

Learning Management System

SUNY Oswego will be working with SLN and ITEC to migrate to Blackboard.  Our first pilot will be in the Spring 2015 semester.

Please check the ANGEL website for dates and times for upcoming workshops.

If you are interested in setting up a community group, please visit the ANGEL web site at www.oswego.edu/academics/angel. Community groups are limited to academic and/or university related activities only.

Professional Development

Spring workshops have completed for the spring semester.   We will be participating in the CELT-sponsored Spring BreakOut sessions during the week of May 28.

If you missed the Cyber Security Awareness trainings offered last fall and in January, you may view a recording of it on our CTS Security site.

There is a new professional development website that includes links to training schedules and registrations for workshops offered by CELT, ORSP, ANGEL, Web Development (Ingeniux) and CTS.  Check it out at www.oswego.edu/workshops.

Self-paced, on-demand tutorials are also available from our iTunes U site.  Tutorials covering Microsoft Word 2010, Excel 2010, PowerPoint 2010 and LakerApps Sites are currently available.

The Skillport e-learning software is available to all faculty, staff and students.  Skillport includes business and technology-related courses (including Office 2010) as well as courses and test preps needed to prepare you for many business and technical certifications.   Get started at the Skillport home page.

Software News

SUNY Oswego faculty and staff may purchase a copy of either MS Office Professional Plus 2010 (PC version) or Office 2011 (Mac version) for use on their personal computers for $9.95 plus tax.  This "Home Use Program" is made available to faculty and staff through the campus purchase of the Microsoft Campus Agreement.  Details of how to go about purchasing can be found from the CTS MS Office Home Use Program (HUP) page.

Microsoft offers students a significantly reduced price on its products.  For the special pricing and to buy online visit www.theultimatesteal.com.

Web Update

Campus Technology Services (CTS) would like to inform users that new releases of browsers aren't always fully compatible with some web-based campus services that you need to access.  The use of an unsupported browser can result in inconsistent or unpredictable behavior when working in applications.  Please check with the TSC before updating.

Desktop Support

During the month of April, Campus Technology Services technicians have continued to aid faculty, staff and students in resolving hardware, software, and account related issues. Since the beginning of April Campus Technology Services technicians have assisted in the resolution of 55 hardware, 77 software, and 18 account related support requests.

Equipment Deployments

Our desktop support technicians continue to prepare and deploy new equipment as
it arrives. During the month of April new equipment has been configured and installed for: Biological Science Department, Campus Center Box Office, Campus Technology Services Department, Environmental Research Center, Facilities Design and
Construction, Graduate Research Office, Interdisciplinary Programs and Activities Center, International Education Department, Laker Box Office, President's Office, and the Recreation Department.

Lab Utilization Statistics

The table below displays the number of login sessions that have occurred within the campus open access computer labs over the last month.

Location # of Logins
Campus Center 207 5975
Hart Lab 919
Mahar 210 863
Mahar 211 1159
Oneida Lab 801
Penfield 24 Hour 3781
Penfield Lab 8356
Penfield Laptops 27
Penfield Lobby 5419
Scales Lab 390
Snygg 6 1040
VDI 1390

During the month of March, 80 virtual applications were launched.

 

Banner

  • Admissions - completed nine program modifications as requested, attended webinar for Common Application supplement developments
  • Advancement/Development - assisted with iModules race/ethnicity problem resolution, revised iModules extract to include student constituents, developed procedures for doing calculations from data extraced from the gift analysis report, reviewed use of ATVRATE codes, trained additional staff on the use of Advancement Self-Service
  • Athletics - assisting with ACS Athletics implementation
  • Extended Learning - developed Veteran Information report, completed one program modification request, and attended meetings to discuss high school program enrollment
  • Financial Aid -  refined batch posting for required tax documents, developed a program to track students that have outstanding charges, assisted in ad hoc reporting as requested; implemented Advanced Queues to manage SaaS Needs Analysis calcs
  • Housing - Assisted Res Life & Housing in resolving issue of our current new student students needing to be processed as a returning student, assisted in explaining a Millennium access problem
  • Institutional Research - SIRIS support,  modified nine programs as requested, developed a program to track students with financial holds that could not pre-register
  • Parking - begin modifications to record individual ticket information in Banner
  • Provost (Extended Learning) - initiated the myOswego veteran information survey
  • Registration - Degreeworks implemenation support, provided Evisions Argos degreeworks reporting as requested, set up web applications for Aspire Higher and consulted on admitting exchange students
  • Student Accounts - began development a new Argos cube as requested; completed data updates as requested to correct voucher errors
  • Student Advisement - Added registered hours to Vizual Zen extract, developed a data extract for Collegiate Link (organization tracking) project
  • Walker Health Center - provided additional immunization data, added active student indicator to daily student extract, and began developing a program to upload Medicat compliance results to Banner
  • Ellucian Live 2013 was attended by Eric Goodnough, Susan Fisk, Lori Reed, and Suzanne Maniccia

Other Administrative Support

  • Banner Upgrade - installed recent Finaid and ResWeb releases on production; Finance, ARSYS, and Alumni releases on test
  • Banner Patches - reviewed list of uninstalled patches to identify which patches needed application after the upgrade (above), installed on test and prod databases
  • Banne XE - configured a new test server to run Banner XE
  • Clean Address Update
  • LDAP Account creation replacement research
  • BDMS support
  •  Created new VM for Tim Yager to upgrade Pharos-srv server
  • Make sure all Faculty/Staff computers get upgrade to Avecto
  • Upgraded Postgres DB server for Theatre Manager DB
  • Upgraded Test Theatre Manager to version 9.21.06
  • Upgraded Test Apache for Theatre Manager ticket website to version 2.4
  • Updated Add to domain application to include 802.1x wired installer.

Provided AV support for the following events

    Non-Campus Events:
  • Lt. Governor's Regional Economic Development Council 
  • Governor's Budget Signing Tour

    Campus sponsored events:   

  • College Council
  • Quest
  • Honors
  • General Faculty Meeting
  • Three Admissions Open Houses
  • School of Communication Media and the Arts Open House
  • Primate Sanctuary Show
  • Joint Chairs Committee Meeting
  • David Benioff Oswego Reading Initiative presentation
  • Summer Projects Planning Meeting

    Department or Unit Sponsored Events:

  • Communications Studies Internships meeting
  • Public Speaking Competition
  • Great Lakes Water Quality presentation
  • Hunger Games

        Campus Projects or Support

  • Installation of A/V equipment and cabling in Shineman
  • Onondaga CATV upgrade planning meeting
  •  Installed equipment to send multiple digital signage channels over TV
  •  Started installation of cabling and control systems in Shineman teaching labs and classrooms
  • Torchlight and Commencement Eve Dinner planning meeting
  • Genius Olympiad planning meeting

        Department Projects or Support

  • SAVAC projector installation consultation
  • Installed display and video switcher in Library conference room
  • Provided training and support for History
  • Installed display and cabling in Admissions reception area simplifying operations

          CTS Projects or Support

  • ATC training video revision
  • Repaired digital signage in Funnelle
  • Planning for CATV extender move out of restricted SAVAC area 
Telecommunications 
  •  VoIP phone installations continue
  • Heat Tickets
  • SEI/ SOE Technology Planning
  • UFPO Dig tickets - locating outside plantcable
  • Moves Add Changes
    • 4 Adds
    • 0 Change
    • 45 Disconnects
    • 2 Moves
  • Monthly Billing

Networking

  • Created DNS records for new alumni imodules services
  • Created DNS records for admissions guidebook services
  • Created DNS records for FYE lakelife services
  • Worked with Judicial Affairs on outstanding DMCA cases
  • Worked on 881 podium router configuration and testing from scratch for Science installations
  • Spoke with Bobby Henderson about UPS issues in Mackin - Lee - Laker
  • Provided wireless access for Admissions, Quest, Sustainability, SAPB, Acron events
  • Troubleshot WRVO FTP lag time issues - turned out to be a flood of remote connections on the remote side
  • Replaced Mackin 2nd floor UPS after determining the power control module was faulty, returned old unit
  • Replaced faulty UPS battery in large Penfield UPS unit, returned old unit
  • Replaced battery modules in existing Snygg UPS units before building goes offline, returned old unit
  • Configured switch ports for small ATA's for elevator phones
  • Troubleshot reported wired network issues on the 7th floor Culkin.  Traced back to a single switch in the switch stack that was crashing and causing loss of connectivity while it rebooted.  Opened TAC case and still working with them to determine exactly what is going on with it.  After 3rd reported outage w temporarily moved all end devices off that switch to another switch in the stack while still working with TAC to determine root cause and a fix
  • Looked into a possible wiring issue with a data jack that was caused link/POE flapping to occur on an interface on the 6th floor Culkin.  Created telecom wiring ticket to investigate
  • Looked into stackwise power issue in Rich after seeing messages show up in the logs.  Working with TAC to determine the root cause and solution
  • Troubleshoot known issue in Building 12 with certain conditions being met that cause a unicast traffic flow issue on an end user switchport.  Working with TAC to determine a root cause, we've tied it back to a pre-existing support case and bug ID, trying to get the fix put into the latest version of firmware we are running on the devices
  • Put together port mapping of Town House data closets and created Visio diagrams
  • Added stack cables to Town House A and G for redundancy
  • Added sheldon-ae-strip and wrvo-strip to StruxureWare
  • Troubleshot a large # of wired ports on 4th floor Hart not working.  Traced back to a student with a faulty patch cable going around and plugging into those ports and having them error-disable
  • Looked into WTOP/WNYO VPN access, modified secrets and got it working again
  • Upgraded NYSERNET Internet2 circuit from 25M to 50M
  • Prepped for emergency broadcast message, went over to UP to connect side-car for push-button notifications and was there in case there were issues sending the message
  • Modified ITEC VPN tunnels to add a specific tunnel for VDI server traffic to a new endpoint at ITEC
  • Added ITEC VDI tunnel to service groups on wan appliance
  • Added ATC podium router addresses to split-tunnel policy for FAC-STAFF on new VPN
  • Looked at why PLAR functionality would not work on the ATA-187's and why you couldn't register the 2nd line on each.  Updated firmware to newer firmware that supports PLAR, and also created a blank SIP Dial-Rule for PLAR functionality from a SIP device
  • Looked into how to activate the 2nd line on a SIP ATA for elevator phones.   ATA must provide CPC disconnect tones in order to work properly, created bugID with TAC and put in feature enhancement request to provide CPC tone feature in next firmware release
  • Modified WRVO firewall and campus firewall to allow WRVN IP traffic to/from WRVO's secure network here on campus for AudioVault access
  • Finished up SEI wireless layout locations, sent to Allen and Ty
  • Helped Dave with PKI certificate creation/validation and adding to WLC controllers
  • Verified UCM MoH file settings and that it will work with multicast properly, documented how to do this
  • Started work to get VOIP test lab up and working again for fax testing, re-connected test publisher and test Unity Connection mailboxes
  • Science building addition switch racking in 2nd, 3rd, 4th floors, un-box horizontal managements and organize, un-box PDU strips.  Need to come up with a different mounting solution for the PDU's
  • Still waiting on power to all data closets in the Science building and data rack and cables to be installed in 1st floor of the addition.  Unable to activate any network connectivity until power is applied to those rooms
  • Fixed Sheldon guest room wired access, policies that had the network removed from the policy
  • Created script for Intermapper to monitor ITEC VDI tunnel traffic from info that Jon gathered, determined it won't monitor exactly the way we need it to.  Still trying to figure out another workaround
  • Spoke with network group at CUNY Graduate Center about their Exinda and our experiences with it
  • XBOX/PS3/Wii/other registrations:  1
  • DMCA take down notifications:  4
  • Abuse Notifications:  2

 

Staff Updates

  • 2 new network staff members started in April - Jonathan Fargo and Derek Rice
  • Crestron P201 programming - Perkins
  • BYOT - From distraction to reaction webinar - Hageny
  • Assisted Listening Implications Workshop -- McQuin

 

 

 


Phone Calls to the Technology Support Center (4/1/13-4/30/13):  

751

 
Call Logging Statistics - 4/1- 4/30/2013

Tickets Logged By Time Number of Tickets
8:00 AM - 11:59 AM 255
12:00 PM - 4:30 PM   279
4:31 PM - 9:00 PM 26
Off Hours 3
Total Tickets Logged 563